OpenStudy (anonymous):

in finding the solutions of differential equations by P, Q, method (dy/dx +Py = Q) , when the integrating factor = e ( to the power , integration of P.dx) is it necessary that Q must be in terms of x variable only? Or can terms with y variable be also present?

OpenStudy (anonymous):

When solving an equation with integrating factors, the ODE must first be in "standard linear form." That requires that P and Q both be functions of the independent (x) variable only. standard form: \[\frac{ dy }{ dx } + p(x)y(x) = q(x)\]

OpenStudy (anonymous):

Thank you for your time. I did not understand 'y(x)" part of your explanation. Will you please explain again? What is ODE? When the differential equation is in linear form with the term dy/dx is it compulsory that P and Q are functions of x ONLY ? Does this mean that when dx/dy term is present,P and Q must be variables of y?

OpenStudy (anonymous):

y(x) means that y is a function of x alone. ODE is 'ordinary differential equation'. Yes, P and Q have to be functions of x only, for such an equation to be a Linear Differential equation. you are right, when it has dx/dy, P and Q must be functions of y alone.

OpenStudy (anonymous):

it can be BERNOULLI'S =N of form \[\frac{ dy }{ dx }+P(x)=Q(x)y ^{n}\] where LHS is a function of x and y